1. A loan of ​$5989 borrowed today is to be repaid in three equal installments due...

1.

A loan of

​$5989

borrowed today is to be repaid in three equal installments due in

two

years​,

three

​years, and

five

​years, respectively. What is the size of the equal installments if money is worth

2.3%

compounded quarterly question markquarterly?The payments are each

​$nothing.

​(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as​ needed.)

Enter your answer in the answer box.

2.

What sum of money will grow to

​$2967.63

in

four

years at

6.6​%

compounded quarterly question markquarterly?The sum of money is

​$nothing.

​(Round to the nearest cent as needed. Round all intermediate values to six decimal places as​ needed.)

Enter your answer in the answer box.
Answer #1

1: first we need to compute the effective rate

EAR = (1+APR/m)^m -1

= (1+ 2.3%/4)^4 -1

=0.02319914

Let the equal amount = A

Present value = CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

5989 = A/(1+0.02319914)^2 + A/(1+0.02319914)^3 + A/ (1+0.02319914)^5

5989 = 0.955168 A + 0.933511A + 0.891660A

Hence

A = 5989/2.780339 = 2154.05

2: Present value =Future value/(1+Rate)^number of periods

= 2967.63/ (1+6.6%/4)^(4*4)

=2283.98